AP4-12A (Statement of income presentation: basic EPS) The following information was taken from Riddell Ltd.s adjusted trial balance as at April 30, 2020 Sales revenue Interest revenue Utilities expense Insurance expense Cost of goods sold Distribution expenses Administrative expenses Depreciation expense Interest expense Income tax expense Dividends declared-Common shares Dividends declared-Preferred shares S1,045,800 7,200 24,000 6,000 396,000 230,250 90,250 81,000 16,500 56,700 15,000 20,000 Required a. Prepare a single-step statement of income for the year ended April 30, 2020 b. Prepare a multi-step statement of income for the year ended April 30, 2020 c. Determine Riddells gross margin percentage for the year d. If Riddell had 35,000 common shares outstanding throughout the year, determine the companys basic earnings per share
Answer #1

Answer

  • Requirement ‘a’ SINGLE STEP

Income Statement

for the year ended Apr 30,2020

Revenues:

Sales Revenue

$1,045,800

Interest Revenue

$7,200

Total Revenues

$1,053,000

Expenses:

Cost of Goods Sold

$396,000

Utilities expenses

$24,000

Insurance expense

$6,000

Distribution expense

$230,250

Administrative expense

$90,250

Depreciation expense

$81,000

Interest expense

$16,500

Income tax expense

$56,700

Total Expenses

$900,700

Net Income

$152,300

  • Requirement ‘b’ MULTI STEP

Income Statement

for the year ended Apr 30,2020

Sales Revenue

$1,045,800

Cost of Goods Sold

$396,000

Gross Profits

$649,800

Operating Expenses:

    Selling Expenses:

Distribution expense

$230,250

    Administrative expenses:

Utilities expenses

$24,000

Insurance expense

$6,000

Administrative expense

$90,250

Depreciation expense

$81,000

Total Operating Expenses

$431,500

Operating Income

$218,300

Other Revenues & (Expenses):

Interest Revenue

$7,200

Interest expense

($16,500)

($9,300)

Net Income before Income tax

$209,000

Income tax expense

$56,700

Net Income

$152,300

  • Requirement ‘c’

A

Gross Profits

$649,800

B

Sales Revenue

$1,045,800

C = (A/B) x 100

Gross Profit %

62.13%

  • Requirement ‘d’

A

Net Income

$152,300

B

Preferred Stock Dividend

$20,000

C = A - B

Net Income for common stockholder

132300

D

Outstanding common stock shares

                    35,000

E = C/D

Basic Earnings per share

$                     3.78